An investor purchases a nine-year, 7% annual coupon payment bond at a price equal to par value. After the bond is purchased and before the first coupon is received, interest rates increase to 8%. The investor sells the bond after five years. Assume that interest rates remain unchanged at 8% over the five-year holding period.

The capital gain/loss per 100 of par value resulting from the sale of the bond at the end of the five-year holding period is closest to a:

1. loss of 8.45

2. loss of 3.31

3. gain of 2.75
Answer #1

Value of bond at the end of 5th year = 100*(1+.08)^5

=100*1.469

=146.93

If the bond would have an interest rate of 7%

= 100*(1+.07)^5

=100*1.40255

=140.2551

It is closest to gain of 2.75